Find the parametric equation of x=a(1+cos theta)


1
Expert's answer
2021-10-14T18:17:57-0400

Here "x=a(1+cos \theta)" is already in parametric form as "x =f(\theta)"

We can add "y=a(1-sin \theta)" to make "y=f(\theta)"

Now, both "x" and "y" are functions of "\theta" .